the ratio of blue marbles to red marbles in a box is 4 to 9. if there are 91 blue marbles and red marbles in the box altogether then how many red marbles are there?
Answer by checkley77(12844)

4x+9x=91
13x=91
x=91/13
x=7
4*7+9*7=91
28+63=91
91=91 Therefore there are 28 blue marbles & 63 red marbles in the box.